# Compositional Reasoning with Transformers, RNNs, and Chain of Thought

### v1

@@ Abstract
We study and compare the expressive power of transformers, RNNs, and transformers with chain of thought tokens on a simple and natural class of problems we term Compositional Reasoning Questions (CRQ). This family captures problems like evaluating Boolean formulas and multi-step word problems. Assuming standard hardness assumptions from circuit complexity and communication complexity, we prove that none of these three architectures is capable of solving CRQs unless some hyperparameter (depth, embedding dimension, and number of chain of thought tokens, respectively) grows with the size of the input. We also provide a construction for each architecture that solves CRQs. For transformers, our construction uses depth that is logarithmic in the problem size. For RNNs, logarithmic embedding dimension is necessary and sufficient, so long as the inputs are provided in a certain order. (Otherwise, a linear dimension is necessary). For transformers with chain of thought, our construction uses $n$ CoT tokens. These results show that, while CRQs are inherently hard, there are several different ways for language models to overcome this hardness. Even for a single class of problems, each architecture has strengths and weaknesses, and none is strictly better than the others.
@@ 1 Introduction
Large language models (Touvron et al., 2023; Anil et al., 2023; Achiam et al., 2023) are increasingly used to perform logical reasoning and other problems that require algorithmic thinking. To understand the power and limitations of these models, it is essential to determine what kinds of computational problems they are capable of solving. To this end, a long line of theoretical work has studied the expressive power of various language modeling architectures using simple tasks like copying strings, recognizing formal languages, and determining if a graph is connected (Jelassi et al., 2024; Sanford et al., 2024b; Strobl et al., 2024). Such studies typically provide constructions or prove impossibility results for a particular architecture or paradigm—such as recurrent neural networks, transformers, and transformers with test-time scaling (“chain of thought”)—following the progress of the field.
@@
In this paper, we take a broader view. We compare the abilities of several different language modeling architectures to solve a class of problems, which we call Compositional Reasoning Questions (CRQs). Our findings prove that there are unavoidable trade-offs between the architectures we consider. No architecture is capable of solving CRQs without some strong dependence on the size of the problem. However, each architecture allows us to limit the use of some computational resource at the expense of the others (see Table 1). Thus, the CRQ task provides a crisp and formal way to characterize the essential differences between these architectures.
Our Compositional Reasoning Questions are inspired by the complex, multi-step questions for which LLMs are often used in practice. For example, consider the question: Amongst the record title winners of each Grand Slam, who is the player with the fewest overall titles? Answering this question requires solving several sub-questions. While some of them can be solved in parallel (i.e., finding the record holders for each Grand Slam), the overall question cannot be answered without first gathering the answers to the sub-questions. It shares this structure with arithmetic and boolean expressions like $(6+2)\cdot(4-5)$ and is also similar to word problems and logic puzzles. Problems like these can be thought of as having an underlying tree structure. The leaves are the possible answers, each non-leaf node represents a sub-question, and the overall answer is the answer to the root node, see Figure 1 for an example. This type of compositional reasoning is a fundamental challenge for LLMs.
@@
Deep transformer (Section 4) $O(\log^{2}n)$ $O(n^{2}\,\log^{2}n)$ $O(\log n)$
RNN (Section 5) $O(\log n)$ $O(n\log n)$ $O(n)$
Chain of Thought (Section 6) $O(1)$ $O(n^{2})$ $O(n)$
@@
    In Section 3.1, we present Compositional Reasoning Questions, a simple, formal framework based on semantic similarity for studying LLM reasoning on arbitrary tree structures.
@@
    In Section 4, we prove that transformers with constant depth cannot solve arbitrary CRQs (Theorem 4.3), but transformers with depth $L$ can solve all CRQs of depth up to $L$ (Theorem 4.1).
@@
    In Section 5, we prove that RNNs with constant hidden dimension cannot solve arbitrary CRQs (Theorem 5.5), but RNNs with $O(\log n)$ hidden dimension and constant depth can solve all CRQs of size $n$ (Theorem 5.4). This ability depends on the inputs being arranged in a particular order (Algorithm 1); if they are ordered adversarially, RNNs require $O(n)$ hidden dimension (Theorem 5.2).
@@
    In Section 6, we prove that transformers augmented with $O(\log n)$ CoT tokens cannot solve CRQs of size $n$, but transformers augmented with $O(n)$ CoT tokens can (Theorem 6.1).
These results demonstrate the trade-offs between different models (Table 1). While deep transformers are highly parallelizable, they require $O(\log n)$ depth in the worst case, so the model size must depend (albeit mildly) on the problem size. Likewise, RNNs use little compute but must grow to handle larger problems. Chain of thought allows a single, constant-size model to handle any CRQ, but it runs slowly and is not parallelizable. Overall, our work reveals a nuanced complexity landscape for an important class of reasoning problems.
@@ 2 Related Work
[…]
Past work has also studied the expressive power of transformers with logarithmic-depth. (Sanford et al., 2024a, b) focus on graph tasks, showing how the parallelism of transformers gives them an advantage over other architectures. Liu et al. (2022) show the ability of logarithmic-depth transformers to simulate finite automata, and Merrill and Sabharwal extend these results by weakening their assumptions. We note that as implied by our Lemma 4.2, CRQs in general cannot be solved by finite-state automata.
@@
Chain of thought (CoT) (Reynolds and McDonell, 2021; Wei et al., 2022b; Nye et al., 2021) is a method that enhances the ability of transformers to solve logical reasoning tasks. Rather than producing the answer all at once, the model is allowed to autoregressively generate a series of intermediate tokens to help it carry out each step in the solution. Li et al. (2024) show that constant-depth transformers that generate $T$ CoT tokens can simulate Boolean circuits of size $T$. Merrill and Sabharwal (2023b) prove that they can simulate Turing machines that run in $O(T)$ steps.
[…]
@@ 3 Problem Formulation and Preliminaries
Notations.
We use bold-face letter to denote vectors, e.g. $\mathbf{x},\mathbf{y}$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and $1\leq i<j\leq d$ we define by $(\mathbf{x})_{i:j}\in\mathbb{R}^{j-i+1}$ the vector whose entries are the entries of $\mathbf{x}$ in places $i$ until $j$. We also define $(\mathbf{x})_{-1}$ the last coordinate of $\mathbf{x}$. For an integer $n$ we denote by $[n]=\{1,\dots,n\}$.
Let $T=(V,E)$ be a rooted tree with root $v_{r}\in V$. The depth of a node $v\in V$ is the length of the unique path from $v\in V$ to $v_{r}$. The depth of the tree is the largest depth of its nodes. The parent of a node $v\in V$ is the node connected to $v$ on the path to the root and denoted by $\mathcal{P}(v)$. The children of $v$ are the nodes whose parent is $v$, and are denoted by $\mathcal{C}(v)$. The degree of a node is the number of its children. Thus, leaves have a degree of 00. We say that a node $u$ is a descendant of $v$ if there is $i$ such that $\mathcal{P}^{(i)}(u)=v$, where $\mathcal{P}^{(i)}$ is the parent function composed with itself $i$ times. We also say in this case that $v$ is an ancestor of $u$.
@@ 3.1 Compositional Reasoning Questions
We formally define the problem of Compositional Reasoning Questions as follows:
@@
The size of the CRQ is defined as $|V|$. The answer to the CRQ is defined as $\mathcal{A}(v_{r})$ where the function $\mathcal{A}$ is defined recursively: For a leaf $u\in V$ we define $\mathcal{A}(u)=\mathbf{x}_{u}$. For a non-leaf node $v\in V$ we define:
@@
In the above definition we assume that the CRQs are defined over some finite vocabulary $\Gamma$. We can think about the vocabulary as $\Gamma=\{0,\pm 1,\dots,\pm 9\}$ for simplicity, although other vocabularies can be used (e.g. $\Gamma=\{\pm 1\})$.
CRQs are thus inherently hierarchical tasks, where the usual sequential structure is replaced by a tree structure. As a result, we will consider learning CRQs with models that can leverage this tree structure. Each node $\mathbf{v}$ is embedded as the vector $(\mathbf{x}_{v},\mathbf{z}_{v},\mathbf{z}_{\mathcal{P}(v)},\ell(v))$, where $\mathbf{x}_{v}$ is the label, $\mathbf{z}_{v}$ is a positional embedding, and $\ell(v)$ is the depth of $v$ in the tree.
@@ 3.2 Transformers
We now provide a formal definition of the transformer architecture that will be used throughout the paper. The input sequence is $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathbb{R}^{d}$. We concatenate Positional Encodings (PE) to the inputs, which have the form $\mathbf{e}_{1},\dots,\mathbf{e}_{n}\in\mathbb{R}^{e}$. The positional encodings cannot depend on the values of the vectors $\mathbf{x}_{i}$, only on their positions (namely, on $i$). The input to the transformer is the sequence $\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\mathbf{x}_{i}\\ \mathbf{e}_{i}\end{pmatrix}\in\mathbb{R}^{d+e}$.
The transformer backbone contains alternating layers of hardmax, single-head self-attention and feed forward networks with skip connections. For simplicity, we do not use layer normalization or attention masking (though the lower bounds generally hold against models that include them). Formally, let $\bm{h}_{i}^{(\ell)}$ denote the hidden embedding output after $\ell$ transformer blocks corresponding to the $i$th input. Define:
@@
where $\bm{V}_{\ell},\bm{K}_{\ell},\bm{Q}_{\ell}\in\mathbb{R}^{(d+e)\times(d+e)}$ and $\theta_{\ell}$ are weight matrices and $\sigma(\cdot)$ is the ReLU function. Ties in the $\operatorname*{arg\,max}$ operation are broken arbitrarily.
Finally, an unembedding layer discards entries corresponding to the positional encoding:
$\mathbf{y}_{i}=\begin{pmatrix}\bm{I}_{d}&\cdot\\ \cdot&\cdot\end{pmatrix}\bm{h}_{i}^{(L)}$ (4)
We consider $\mathbf{y}_{n}$ to be the output of the model. Following previous work (Merrill and Sabharwal, 2023a, c), we assume that the inputs, weights, and intermediate representations of the network can all be represented using $O(\log n)$ bits of precision.
@@ 4.1 Deep transformers can solve CRQs
This subsection presents our construction of a deep transformer that solves CRQs:
@@
For any $L,n\in\mathbb{N}$ there exists a transformer $T$ with depth $L-1$ and embedding dimension $O(d+\log(n))$ that solves all the compositional reasoning questions up to $n$ nodes and depth $L$.
The proof appears in Section A.1. First, note that the depth of the transformer depends only on the depth of the tree of the CRQ. This is one of the main strengths of transformers: parallelism. Namely, the transformer is able to solve all the sub-questions in the tree of the same depth with just one layer.
@@
We now give a short intuition for our construction. We first define a positional encoding vector for each node that has four parts. For each non-leaf node $v$, we define an identifying vector $\mathbf{z}_{v}\in\{\pm 1\}^{O(\log(n))}$ such that $\left\langle\mathbf{z}_{v},\mathbf{z}_{u}\right\rangle\ll\left\lVert\mathbf{z}_{v}\right\rVert^{2}$ for any nodes $u,v\in V$. For leaf nodes, we let $\mathbf{z}_{v}=\boldsymbol{0}_{O(\log(n))}$. This is the first part. The second part of node $v$’s positional encoding vector is the identifier of its parent, $\mathbf{z}_{\mathcal{P}(v)}$. The third part is the depth of $v$ in the tree. The fourth is an indicator variable that is $1$ when the depth is $L-1$, where $L$ is the depth of the tree, and 00 otherwise. These four parts are all concatenated to the value of the node $\mathbf{x}_{v}$.
@@
  • Each non-leaf tokens that have depth smaller than $L-1$ attends only to itself.
  • Each non-leaf tokens that has depth $L-1$ attends only to its children. Specifically, it attends to the child child that answers its sub-question, one whose value vector has the largest correlation with the non-leaf token.
@@
By constructing this attention pattern, all the sub-questions in layer $L-1$ are solved simultaneously. We then use the MLP to reorganize the tokens. We increase the counter of the depth of every token, and update the indicator for nodes of depth $L-1$. Nodes that in the previous iteration had depth $L-1$ are altered to resemble leaves, i.e. their $\mathbf{z}_{v}$ vector is set to zero. For these nodes, we also set their value vector $\mathbf{x}_{i}$ to be the answer to their sub-question, as computing in the previous attention layer. In the next attention layer, nodes previous in layer $L-1$ will play the part of leaves and be attended to by their parents, which were previously in layer $L-2$. Applying this construction $L-1$ times will compute the answer to the root, which is the answer to the CRQ.

### v2

@@ Abstract
It is well understood that different neural network architectures are suited to different tasks, but is there always a single best architecture for a given task? We compare the expressive power of transformers, RNNs, and transformers with chain of thought tokens on a simple and natural class of tasks we term Compositional Reasoning Questions (CRQ). This family captures multi-step problems with tree-like compositional structure, such as evaluating Boolean formulas. We prove that under standard hardness assumptions, none of these three architectures is capable of solving CRQs unless some hyperparameter (depth, embedding dimension, and number of chain of thought tokens, respectively) grows with the size of the input. We then provide constructions for solving CRQs with each architecture. For transformers, our construction uses depth that is logarithmic in the problem size. For RNNs, logarithmic embedding dimension is necessary and sufficient, so long as the inputs are provided in a certain order. For transformers with chain of thought, our construction uses $n$ CoT tokens for input size $n$. These results show that, while CRQs are inherently hard, there are several different ways for language models to overcome this hardness. Even for a single class of problems, each architecture has strengths and weaknesses, and none is strictly better than the others.
@@ 1 Introduction
Large language models (Touvron et al., 2023; Anil et al., 2023; Achiam et al., 2023) are increasingly used to perform logical reasoning and other problems that require algorithmic thinking. To understand the power and limitations of these models, it is essential to determine what kinds of computational problems they are capable of solving. To this end, a long line of theoretical work has studied the expressive power of various language modeling architectures using simple tasks like copying strings, recognizing formal languages, and determining if a graph is connected (Jelassi et al., 2024; Sanford et al., 2024b; Strobl et al., 2024). Such studies typically provide constructions or prove impossibility results for a particular architecture or paradigm—such as recurrent neural networks, transformers, and transformers with test-time scaling (“chain of thought”)—following the progress of the field. [^1] In this paper, we consider transformers with and without chain of thought to be distinct architectures. We also distinguish between “shallow” (constant number of layers compared to the input size) and “deep” transformers. While these are all types of transformer, our results show that they are qualitatively very different.
@@
In this paper, we take a broader view. We compare the abilities of several different language modeling architectures to solve a class of problems that we call Compositional Reasoning Questions (CRQs). While it is intuitive that each architecture has strengths and weaknesses, a rigorous understanding of this phenomenon is incomplete. Theoretical comparisons between architectures often use tasks that are selected to make one model look good. For instance, copying an arbitrarily long string is impossible for an RNN with fixed memory, but easy for a transformer (Jelassi et al., 2024). In this work, we prove that the relationships between model classes can be much subtler and more interesting. Our study of CRQs prove that there are unavoidable trade-offs between the architectures we consider, even for a single problem. None of these architectures can solve CRQs without a strong dependence on the size of the input. However, each architecture allows us to minimize the use of some computational resource compared to the others (see Table˜1). CRQs are simultaneously hard in different ways for different architectures, but not intractable for any one of them. Thus, the CRQ task provides a crisp and formal way to characterize the essential differences between these architectures.
Our motivation for defining and studying Compositional Reasoning Questions is as follows. Many multi-step reasoning tasks share a common tree-like structure, meaning that they are hierarchical compositions of smaller tasks. Consider the following examples:
  • What is $(6+2)\cdot(4-5)$?
  • Among all U.S. states, which one’s highest mountain has the lowest elevation?
To answer these questions, we must first solve several sub-tasks, like “What is 6+26+2?” and “What is the elevation of the highest point in Hawaii?” While some of these sub-tasks can be solved in parallel, the overall task cannot be solved without first gathering the answers to the sub-tasks. Furthermore, sub-tasks may be nested hierarchically to form more complex tree structures, as in $5\cdot(8-(9\div(2-1)))$. Each sub-task corresponds to a non-leaf node of the tree, and the answer to the overall questions corresponds to the root node. See Figure˜1 for an example tree. The paradigmatic example of compositional reasoning is Boolean formula evaluation, which is the problem of determining the truth-value of an expression like $(\mathtt{T}\lor\mathtt{F})\land(\neg(\mathtt{F}\land\mathtt{T}))$. This is a problem of fundamental importance in basic logic and in complexity theory, where it is one of the key $NC^{1}$-complete problems, but it has been surprisingly understudied in prior work on LLM reasoning. Various other tasks studied in the literature also share this hierarchical structure (Sinha et al., 2019; Feng et al., 2024). Compositional Reasoning Questions, which we define in Section˜3.1, provide a simple and unified framework for studying reasoning tasks that exhibit compositionality. In particular, we prove that the CRQ task captures Boolean formula evaluation. Therefore, we believe that CRQs are of fundamental importance as a yardstick for LLM reasoning.
@@
[…]
@@
    In Section˜3.1, we present Compositional Reasoning Questions, a simple, formal framework based on semantic similarity for studying LLM reasoning on arbitrary tree structures.
@@
    In Section˜4, we prove that transformers with constant depth cannot solve arbitrary CRQs (Theorem˜4.3), but transformers with depth $L$ can solve all CRQs of depth up to $L$ (Theorem˜4.1).
@@
    In Section˜5, we prove that RNNs with constant hidden dimension cannot solve arbitrary CRQs (Theorem˜5.5), but RNNs with $O(\log n)$ hidden dimension and constant depth can solve all CRQs of size $n$ (Theorem˜5.4). This ability depends on the inputs being arranged in a particular order (Algorithm˜1); if they are ordered adversarially, RNNs require $O(n)$ hidden dimension (Theorem˜5.2).
@@
    In Section˜6, we prove that transformers augmented with $O(\log n)$ CoT tokens cannot solve CRQs of size $n$, but transformers augmented with $O(n)$ CoT tokens can (Theorem˜6.1).
Each of these results fills a gap in the literature; taken together, they demonstrate the fundamental trade-offs between different models (Table˜1). While deep transformers are highly parallelizable, they require $O(\log n)$ depth in the worst case, so the model size must depend (albeit mildly) on the problem size. Likewise, RNNs use little compute but must grow to handle larger problems, although their success depends on the order of the inputs. Chain of thought allows a single, logarithmic-size model to handle any CRQ, but it runs slowly and is not parallelizable. Overall, our work reveals a rich complexity landscape for an important class of reasoning problems.
@@ 2 Related Work
Our work belongs to a large body of research studying the representational capacity of transformers, chain of thought prompting and RNNs. Transformers with unbounded depth are known to be universal approximators (Yun et al., 2019), and can simulate Turing machines (Wei et al., 2022a; Merrill and Sabharwal, 2023b) if their size can grow with the sequence length. Looped transformers, which repeat a fixed sequence of transformer layers a variable number of times depending on the length of the input, can implement a simple but universal programming language (Giannou et al., 2023). Several works studied the expressive power of shallow transformers in solving certain representative tasks. Sanford et al. (2024c) study problems like sparse averaging and matching, Yehudai et al. (2024) study counting, and Amsel et al. (2025) study nearest neighbor. A long line of work (Hahn, 2020; Hao et al., 2022; Merrill et al., 2022; Strobl et al., 2024) has used circuit complexity classes to characterize the power and limitations of transformers. Merrill and Sabharwal (2023c) show that constant-depth transformers can only compute functions that are computable by Boolean circuits in the complexity class ${\mathsf{TC}}^{0}$. We use this result to prove that they cannot solve CRQs either (Theorem˜4.3). Finally, Chen et al. (2024) study decoder-only transformers, in which each token can attend only to the previous tokens in the sequence, and prove that $L$-layer decoders are strictly weaker than $L+1$-layer decoders for the task of function composition.
Past work has also studied the expressive power of transformers with logarithmic-depth. (Sanford et al., 2024a, b) focus on graph tasks, showing how the parallelism of transformers gives them an advantage over other architectures. Liu et al. (2022) show the ability of logarithmic-depth transformers to simulate finite automata, and Merrill and Sabharwal (2025) extend these results by weakening their assumptions. We note that as implied by our Lemma˜4.2, CRQs in general cannot be solved by finite-state automata.
@@
Chain of thought (CoT) (Reynolds and McDonell, 2021; Wei et al., 2022b; Nye et al., 2021) is a method that enhances the ability of transformers to solve logical reasoning tasks. Rather than producing the answer all at once, the model is allowed to autoregressively generate a series of intermediate tokens to help it carry out each step in the solution. Li et al. (2024) show that constant-depth transformers that generate $T$ CoT tokens can simulate Boolean circuits of size $T$. Merrill and Sabharwal (2023b) prove they can simulate $O(T)$ steps of a Turing machine.
Most closely related to our work is that of Feng et al. (2024), which studies the power of chain of thought in solving arithmetic problems. Like CRQs, arithmetic expressions correspond to tree of subproblems and are solved by working up the tree. Unlike their work, in which the tree structure is indicated by parentheses in the sequence, we encode the tree structure into the positional encodings of the tokens. However, both problems are ${\mathsf{NC}}^{1}$-hard. Like Feng et al. (2024), we prove that our task is solvable by a constant depth transformer using chain of thought, but not otherwise. However, there are several differences: (i) We provide an explicit construction for solving CRQs using logarithmic depth transformers without chain of thought. (ii) While their task is unsolvable by RNNs with a hidden dimension of $o(n/\log n)$, we provide an explicit construction of an RNN that solves CRQs using only $O(\log n)$, but does not work for all orderings of the input. (iii) Our chain of thought solution generates $n$ tokens, while theirs requires $\Omega(n^{2})$ tokens. Such different conclusions emphasize the importance of the input format; in particular our more favorable scalings hint at the importance of leveraging prior information about the hierarchical structure. In Appendix˜F we generalize our definition of CRQ to also capture arithmetic operations, as studied in Feng et al. (2024).
@@ 3 Problem Formulation and Preliminaries
Notation.
We use bold letters for vectors, e.g. $\mathbf{x},\mathbf{y}$. Let $(\mathbf{x})_{i:j}\in\mathbb{R}^{j-i+1}$ denote entries $i$ through $j$ of $\mathbf{x}$, and $(\mathbf{x})_{-1}$ its last entry. For $n\in\mathbb{N},[n]=\{1,\dots,n\}$. Let $T=(V,E)$ be a rooted tree with root $v_{r}\in V$. The depth of a node $v\in V$ is the length of the unique path from $v\in V$ to $v_{r}$. The depth of the tree is the largest depth of its nodes. The parent of a node $v\in V$ is the node connected to $v$ on the path to the root and denoted by $\mathcal{P}(v)$. The children of $v$ are the nodes whose parent is $v$, and are denoted by $\mathcal{C}(v)$. The degree of a node is the number of its children. Thus, leaves have a degree of 0. We say that a node $u$ is a descendant of $v$ if there is $i$ such that $\mathcal{P}^{(i)}(u)=v$, where $\mathcal{P}^{(i)}$ is the parent function composed with itself $i$ times. We also say in this case that $v$ is an ancestor of $u$.
@@ 3.1 Compositional Reasoning Questions
We formally define the class of Compositional Reasoning Questions as follows:
@@
The size of the CRQ is defined as $|V|$. The answer to the CRQ, denoted as $\mathcal{A}(v_{r})$, is defined recursively: For a leaf $u\in V$ we define $\mathcal{A}(u)=\mathbf{x}_{u}$. For a non-leaf node $v\in V$ we define:
@@
In Section˜A.2, we prove that Boolean formula evaluation is a special case of CRQ. Intuitively, the leaves represent Boolean literals (true and false), the sub-questions represent logical operations (and, or, etc.), the tree topology encodes the order of operations, and the answer is the truth value of the overall expression. Using vector representations is not only more general, but also makes for simple constructions using language modeling architectures like attention.
In the above definition we assume that the CRQs are defined over some finite vocabulary $\Gamma$. We can think of it as $\Gamma=\{0,\pm 1,\dots,\pm 9\}$ for simplicity, although any other vocabulary is allowed too.
CRQs are thus inherently hierarchical tasks, where the usual sequential structure is replaced by a tree structure. Therefore, we will consider learning CRQs with models that can leverage this tree structure. Each node $\mathbf{v}$ is embedded as the vector $(\mathbf{x}_{v},\mathbf{z}_{v},\mathbf{z}_{\mathcal{P}(v)},\ell(v))$, where $\mathbf{x}_{v}$ is the label, $\mathbf{z}_{v}$ is a positional embedding, and $\ell(v)$ is the depth of $v$ in the tree.
@@ 3.2 Transformers
We now provide a formal definition of the transformer architecture that will be used throughout the paper. The input sequence is $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathbb{R}^{d}$. We concatenate Positional Encodings (PE) $\mathbf{e}_{1},\dots,\mathbf{e}_{n}\in\mathbb{R}^{e}$ to the inputs. The positional encodings cannot depend on the values of the vectors $\mathbf{x}_{i}$, only on their positions (namely, on $i$). The input to the transformer is the sequence $\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\mathbf{x}_{i}\\ \mathbf{e}_{i}\end{pmatrix}\in\mathbb{R}^{d+e}$. The transformer backbone contains alternating layers of hardmax, single-head self-attention and feed forward networks with skip connections. For simplicity, we do not use layer normalization or attention masking (though the lower bounds like Theorem˜4.3 generally hold against models that include them). Formally, let $\bm{h}_{i}^{(\ell)}$ denote the hidden embedding output after $\ell$ transformer blocks corresponding to the $i$th input. Define:
@@
where $\bm{V}_{\ell},\bm{K}_{\ell},\bm{Q}_{\ell}\in\mathbb{R}^{(d+e)\times(d+e)}$ and $\theta_{\ell}$ are weight matrices and $\sigma(\cdot)$ is the ReLU function. Ties in the $\operatorname*{arg\,max}$ are broken arbitrarily. Finally, an unembedding layer $\mathbf{y}_{i}=\begin{pmatrix}\bm{I}_{d}&\cdot\\ \cdot&\cdot\end{pmatrix}\bm{h}_{i}^{(L)}$ discards entries corresponding to the positional encoding. We consider $\mathbf{y}_{n}$ to be the output of the model. Following previous work (Merrill and Sabharwal, 2023a, c), we assume that the inputs, weights, and intermediate representations of the network can all be represented using $O(\log n)$ bits of precision. We use hardmax attention for ease of analysis, following previous works on the theory of transformers. However, all our constructions can be extended to softmax. To do so, we would modify the statements of Theorems˜4.1 and 6.1 to say that our constructions approximate the target function up to arbitrary error $\epsilon>0$ that is independent of $n$, rather than exactly expressing the target. The proofs would then tune the temperature of the softmax function so that it is sufficiently close to hardmax to achieve approximation error $\epsilon$.
@@ 4.1 Deep transformers can solve CRQs
For any $L,n\in\mathbb{N}$ there exists a transformer $T$ with depth L1L-1 and embedding dimension $O(d+\log(n))$ that solves all CRQs with at most $n$ nodes and depth at most $L$.
The proof appears in Section˜A.1. First, note that the depth of the transformer depends only on the depth of the tree of the CRQ. This is one of the main strengths of transformers: parallelism. Namely, the transformer is able to solve all the sub-questions in the tree of the same depth with just one layer.
@@
We now give a short intuition for our construction. We first define a positional encoding vector for each node that has four parts. For each non-leaf node $v$, we define an identifying vector $\mathbf{z}_{v}\in\{\pm 1\}^{O(\log(n))}$ such that $\left\langle\mathbf{z}_{v},\mathbf{z}_{u}\right\rangle\ll\left\lVert\mathbf{z}_{v}\right\rVert^{2}$ for any nodes $u,v\in V$. For leaf nodes, we let $\mathbf{z}_{v}=\boldsymbol{0}_{O(\log(n))}$. This is the first part. The second part of node $v$’s positional encoding vector is the identifier of its parent, $\mathbf{z}_{\mathcal{P}(v)}$. The third part is the depth of $v$ in the tree. The fourth is an indicator variable that is $1$ when the depth is L1L-1, where $L$ is the depth of the tree, and 0 otherwise. These four parts are all concatenated to the value of the node $\mathbf{x}_{v}$.
@@
  • Each non-leaf tokens that have depth smaller than L1L-1 attends only to itself.
  • Each non-leaf tokens that has depth L1L-1 attends only to its children. Specifically, it attends to the child child that answers its sub-question, one whose value vector has the largest correlation with the non-leaf token.
@@
By constructing this attention pattern, all the sub-questions in layer L1L-1 are solved simultaneously. We then use the MLP to reorganize the tokens. We increase the counter of the depth of every token, and update the indicator for nodes of depth L1L-1. Nodes that in the previous iteration had depth L1L-1 are altered to resemble leaves, i.e. their $\mathbf{z}_{v}$ vector is set to zero. For these nodes, we also set their value vector $\mathbf{x}_{i}$ to be the answer to their sub-question, as computing in the previous attention layer. In the next attention layer, nodes previous in layer L1L-1 will play the part of leaves and be attended to by their parents, which were previously in layer L2L-2. Applying this construction L1L-1 times will compute the answer to the root, which is the answer to the CRQ.